OpenStudy (anonymous):

Which polynomial, written in simplified form, represents the area of this rectangle?

OpenStudy (anonymous):

a. 8x2-36xy-20y2 b. 8x2+22xy-20y2 c. 16x2+72xy-40y2 d. 8x2+36xy-20y2 Please show your work on how you got this. By the way here is the rectangle

OpenStudy (anonymous):

Area of rectangle = length * breadth/width here length = 8x - 4y and breadth = x + 5y so area of rectangle = (8x - 4y)(x + 5y) = 8x(x + 5y) - 4y(x + 5y) = 8x^2 + 40xy - 4xy - 20y^2 = 8x^2 + 36xy - 20y^2 so d) is the right answer